OpenStudy (rainbow_dash):

An armadillo has a litter of 4. What is the probability that all are 4 males?

OpenStudy (anonymous):

4/4 = 1

OpenStudy (anonymous):

it equal 1

OpenStudy (rainbow_dash):

choices are 1/8, 1/16, 1/32, 1/256

OpenStudy (anonymous):

its 1/16

OpenStudy (anonymous):

Because the probability for each individual being male is 1/2 1/2*1/2*1/2*1/2 =1/16 OK?

OpenStudy (anonymous):

yep

OpenStudy (rainbow_dash):

okay sounds good :)